OpenStudy (anonymous):

2. How many real number solutions are there to the equation 0 = –3x² + x – 4? (1 point) A. 0 B.1 C.2 D.3

OpenStudy (kinggeorge):

First step, calculate the discriminant. So the discriminant is \[1^2-4(-3)(-4)=1-48=-47\]Since this is negative, it must contain zero real solutions.

OpenStudy (anonymous):

so A is correct?

OpenStudy (kinggeorge):

Yes.

OpenStudy (anonymous):

thanks

OpenStudy (kinggeorge):

you're welcome.